# Machine Learning-Based Ultrasonic Weld Characterization Using Hierarchical Wave Modeling and Diffusion-Driven Distribution Alignment

### v1

@@ Abstract
Automated ultrasonic weld inspection remains a significant challenge in the nondestructive evaluation (NDE) community to factors such as limited training data (due to the complexity of curating experimental specimens or high-fidelity simulations) and environmental volatility of many industrial settings (resulting in the corruption of on-the-fly measurements). Thus, an end-to-end machine learning (ML) workflow for acoustic weld inspection in realistic (i.e., industrial) settings has remained an elusive goal. This work addresses the challenges of data curation and signal corruption by proposing workflow consisting of a reduced-order modeling scheme, diffusion based distribution alignment, and U-Net-based segmentation and inversion. A reduced-order Helmholtz model based on Lamb wave theory is used to generate a comprehensive dataset over varying weld heterogeneity and crack defects. The relatively inexpensive low-order solutions provide a robust training dateset for inversion models which are refined through a transfer learning stage using a limited set of full 3D elastodynamic simulations. To handle out-of-distribution (OOD) real-world measurements with varying and unpredictable noise distributions, i.e., Laser Doppler Vibrometry scans, guided diffusion produces in-distribution representations of OOD experimental LDV scans which are subsequently processed by the inversion models. This integrated framework provides an end-to-end solution for automated weld inspection on real data.
@@ 2   Acoustic Model and Forward Problem
Curating a sufficiently large and diverse datasets that faithfully represent the dominant characteristics of the forward problem (i.e., steady-state wave propagation over weldlines) is a necessary precursor for any ML-based inversion task. This is a challenge for weld inspection [38] since high-fidelity simulations capturing all conceivable phenomena (e.g., anisotrpy, nonlinearity, etc) are notoriously expensive and challenging to perform [39]. While some researchers have utilized generative techniques to address this [27, 30], such methods still require a large dataset to train augmentative models or risk biasing generations toward a narrow training distribution, limiting dataset diversity. [40]. Hence, we propose a hierarchy of forward model complexities as an alternative. First, a high-fidelity 3D elastodynamic Navier-Lamé (NL) model that captures the dominant scattering physics, but at a high computational expense. Next, a low-order 2D effective medium (EM) model utilizing Lamb wave theory to construct 2D scalar-valued models that capture the dominant wave behavior of the 3D elastodynamic process at a reduced computational cost. The NL model, being far more expensive to query, serves both to calibrate EM solutions and to provide a sparse datasets (roughly 10% of the entire training set) of full fidelity solutions for ML model tuning.
@@ 2.1   Elastodynamic NL Model
The NL model is constructed as follows. Let $\Omega=[0,L]\times[0,W]\times[0,H]\in\mathbb{R}^{3}$ with external boundary $\partial\Omega$ be the 3D scattering domain with L,W,H>0L,\ W,\ H>0 being the length, width, and height. Let $\Omega_{w}\subset\Omega$ be a welded region along a 2D path $\gamma(s),s=(x(s),y(s))$ with narrow voids representative of cracks in weldlines, $\Gamma_{c}\subset\Omega_{w}$. In some simulations, perfectly matched layers (PMLs) are introduced along $\partial\Omega$, denoted as $\Omega_{\mathsf{PML}}$. Time-harmonic tractions are applied in $\Gamma_{t}\subset\partial\Omega$ denoted as $\bm{f}(\bm{x})=-F(x,y)\exp(i\omega t)\bm{e}_{z}$, with $F(x,y)$ being the amplitude profile on the surface $z=h=H/2$, with $\bm{x}=\begin{bmatrix}x&y&z\end{bmatrix}^{\intercal}$ the being position vector, $\omega$ the frequency, and $i$ the imaginary unit. Defining $\Omega_{b}:=\Omega\setminus\Omega_{w}$ as the bulk material region, the interface $\Gamma_{wb}:=\partial\Omega_{w}\cap\partial\Omega_{b}$ is taken as the boundary between weld and bulk-material such that $\overline{\Omega_{w}}\cup\overline{\Omega_{b}}=\overline{\Omega}$. We consider constant material properties within $\Omega_{b}$ and spatially varying properties in $\Omega_{w}$.
@@
The spatial variations of $E(\bm{x})$ are designed to emulate weldline behavior per the baseline material stiffness $E_{0}$. Namely, we consider a composition of several sub-functions: a nominal weldline function $W(\bm{x})$, a bead function ${B}(\bm{x})$ and a pseudo-randomly structured variation function ${V}(\bm{x})$. $W(\bm{x})$ enforces a baseline stiffness reduction (common in weldlines) about a weld radius $r_{w}$, ${B}(\bm{x})$ applies a pseudo-periodic breadline profile to stiffness reduction line (Fig 1(a)), and ${V}(\bm{x})$ takes the form of a smoothly varying Gaussian random field so as to emulate heterogeneity in weldline strength. Altogether, $E(\bm{x})$ is given as,
@@
For NL simulations, we utilize quadratic Lagrange tetrahedral elements using the gmsh-4.3.1 API for meshing, FEnICsx-0.6.0 for system matrix assembly, and PETSc-3.18.5 as the solver engine. Further details on the FE implementation of Eq (1) are given in B. Even with efficient numerical solvers, there is a high memory requirement associated with the 3D vector-valued NL problem on a high-fidelity mesh (Fig 1(c)). This is particularly the case for (relatively) high frequency simulations, whereby at least 6 elements per wavelength are required for quadratic elements; this is further exasperated by the complexity of welded domains as sufficient mesh refinement is required near the weld line and crack to account for the rapidly varying material properties. For example, a baseline simulation with W=L=12W=L=12", H=0.25H=0.25", and $\omega=2\pi\times 225000$ rad/s necessitates millions of degrees of freedom. Therefore, a lighter simulation procedure is desired; we describe a proposed alternative the subsequent section.
@@ 2.2   Effective Medium Model
Figure 3: Confirmation of the Lamb-wave assumptions. (a) A full D continuum for computing NL solutions and reduced 2D continuum for generating EM solutions. (b) The resulting out-of-plane displacement fields at $z=h$ with their filtered (c) $(A,0)$ and (d) $(S,0)$ components showing qualitative agreements. (e) The spectra of each wavefield show aggregated energy along pre-computed Lambe wave mode numbers (dashed-lines), confirming the alignment of the high-fidelity problem with Lamb theory.
@@ 2.5   Dataset Design
Simulation datasets were generated to ensure sufficient diversity of key domain parameters in both training and testing. Namely, the weld width $L_{w}$, weld angle $\theta_{w}$, force location $\bm{x}_{f}$, nominal weld stiffness reduction $W_{0}$, and relative weld variation amplitude $V_{0}$, are drawn form normal distributions—this emulates deviations from a nominal welding and inspection processes due to natural process variability from nominal welding and measurement schemes. The crack length $L_{c}$ and crack depth $d_{c}$ are both drawn from uniform distributions—this emulated deviations in crack nature which are not suspected to be biased toward any nominal length or depth. Finally, we consider problem domains both with and without cracks, with the probability of a crack occurring denoted as $\mathcal{P}_{c}$ begin sampled from a Bernoulli—this prevents biasing the proposed data-driven inversion framework from always predicting cracks when none are present. These parameters are varied per the distributions defined in Table 3 for each problem domain discussed in section 2.4 with the exception of $\theta_{w}$ which is fixed at 0 for the periodic pipe problem to ensure continuity for the periodic boundary conditions. The distributions of Table 3 ensure that the models are trained on a diversity of meta-parameters emulating real-world variations. For instance, scanning fixtures such as scanning laser doppler vibrometers (LDVs) may not always be perfectly oriented, and hence the weld may appear shifted or tilted relative to the scan. Moreover, the placement of transducers is subject to change, so $\bm{x}_{f}$ in the simulated models should emulate this. Lastly, weldlines may vary in width and relative strength, and these nominal parameters must be diversely sampled as well to mitigate bias in training.
@@ 3   Machine Learning Framework
The machine learning methodology addressed two tasks: wavefield inversion and distribution alignment. The former is a set of inverse models that are responsible for ingesting waveifeld data, either directly from simulation test sets, real-world experiments, or from the output of a conditional generation. The latter serves to draw out-of-distribution (OOD) experimental samples into the training distribution, i.e., the EM and NL simulations, via conditional generative modeling.
@@ 3.1   Inversion Models
The baseline architectures of $\mathsf{U}_{{\text{inv}}}$ and $\mathsf{U}_{{\text{seg}}}$ are both comprised of 5 encoding and decoding blocks, with convolution block attention (CBAM) layers begin applied after each convolution block. Aside from this, $\mathsf{U}_{{\text{seg}}}$ incorporates batch normalization to stabilize training, as well as a multi-axis feature extraction (MAFE) module following CBAM to improve spatial contextualizing in the segmentation task, whereas $\mathsf{U}_{{\text{inv}}}$ incorporates Fourier Neural Operator (FNO) modules between skipped connections to capture global dependencies efficiently in the frequency domain, facilitating accurate modeling of complex acoustical fields. The workflow draws inspiration form the recent results of Ref [43], which demonstrated the advantages of FNO, CBAM, and MAFE modules for forward modeling of discontinuous parametric domains. The details of each module are detailed below.
@@ 3.1.3 Multi-Axis Feature Extraction
The Multi-Axis Feature Extraction (MAFE) module is designed to capture both global and local dependencies along spatial axes in order to strengthen contextual representation for segmentation; similar methodologies has proven effective in various image segmetnation tasks [48, 49]. Herein, we utilize the same MAFE module described Ref [43]. Given an input feature map ${\mathsf{F}}\in\mathbb{R}^{B\times C\times H\times W}$, where $B$, $C$, $H$, and $W$ denote batch, channel, height, and width, respectively, MAFE applies axis-wise gated MLPs along the horizontal ($f_{m_{x}}$) and vertical ($f_{m_{i}}$) directions after normalization, producing a global response, $\mathsf{g}=f_{m_{x}}({\mathsf{F}})+f_{m_{y}}({\mathsf{F}})$. To complement this with local information, the feature map is partitioned into non-overlapping blocks of size $\tfrac{H}{G}\times\tfrac{W}{G}$ ($G=2$ herein), unfolded into patches, and processed by local axis MLPs $f^{l}_{x},f^{l}y$. This yields the local response
@@ 3.1.4 Inverse Model Training Objectives
Three competing objectives guide the weld parameter regression model. The first is an MSE loss between the true and predicted paramter map on $\bar{\bm{x}}$, denoted $\phi(\bar{\bm{x}})$ and scaled by $\mathsf{w}_{1}$—it purpose is self-evident. Second, a gradient loss encourage smoothness in non-welded (flat) regions, which is scaled by $\mathsf{w}_{2}$. Finally, a focal loss scaled by $\mathsf{w}_{3}$ exploits local and spatially compact variations. The composite loss is thus,
@@ 3.2   Super-resolution Conditional Diffusion Model
The inversion framework outlines in section 3.1 is exclusively trained on simulated datasets. Ergo, generalizing its performance in a zero-shot fashion to experimental measurements—for which data is too scarce to train with—must be achieved. Experimental LDV meta-parameters alter the underling distribution of noise and artifacts in unpredictable ways, making it difficult to conventionally denoise experimental measurements or to synthetically noise training data to represent the diversity of noise distributions that may be present in live deployment. Moreover, even the high-fidelity NL solutions are simplified in the sense that they neglect nonlinearity and anisotrpy, an thus miss the full complexity of real-world measurements, meaning that even clean signals require distribution re-alignment when processing experimental data. Therefore, we propose an alternative to conventional data pre-processing—a generative process that is trained to produce wavefield representations from the training (simulation) distribution based on sparse or corrupted wavefield representations from physical measurements. Specially, we consider a conditioned denoising diffusion probabilistic model (DDPM), detailed herein.
@@ 4   Results
We utilize numerical and physical experiments to evaluate the enrichment training scheme, the model performance on hold-out NL simulations, and the adaptability to real-world data. We first demonstrate the Em solutions (2.2) provide meaningful data enrichment for learning the inverse of high-fidelity NL models (2.1).This is followed by qualitative and quantitative performance metrics for each of the problem classed described in section 2.4. Lastly, the performance of the DDPM-driven distribution alignment described by section 3.2 and downstream inverse task is shown for the experimental data described in section 2.6.
@@ 4.2   Experimental Results
The inference results on direct measurements are given by Fig 13(a), showing the real-component of $\bm{\varphi}^{\text{exp}}$ for each scan speed as well as the model’s crack and weld stiffness predictions on $\mathsf{X}^{\text{exp}}$. The models struggle to achieve either weld characterization or crack identification on the direct measurements. Although the large crack of the experimental sample is identified for all scan speeds, it is accompanied by many false positives. As the scan rate increases, the smaller crack is lost, and the predicted crack maps seem to loose all meaning with respect to the ground truth depicted in Fig 5. Moreover, the parameter maps produced by $\mathsf{U}_{\rm inv}$ are difficult to interpret, with discontinuities and a non-physically noisy boundaries shown. Hence, as expected, the model fails to generalize to OOD experimental measurements directly.
Fig 13(b) depicts the inverse same results applied to conditionally generated fields, $\bm{\varphi}^{\text{exp}}_{0}$. Now, in contrast, exceptional performance is recovered for the crack detection problem, with the predicted mask $\hat{\mathcal{C}}\left(\mathsf{X}^{\text{exp}}_{0}\right)$ indicating cracks in the same two regions flagged by the calibrated ground-truth mask of Fig 5. Moreover, the inference for the weld properties are now consistent with the expected result: nominal bulk properties in the unwelded domain, and a noticeable (15-20 percent) stiffness reduction with in the weldline itself. Hence, the proposed framework distribution of OOD experimental measurements to the known training distribution (based on 3D elastodynamic modeling) is shown effective for applying the inverse model to real-world measurements.

### v2

@@ Abstract
Automated ultrasonic weld inspection remains a significant challenge in the nondestructive evaluation (NDE) community to factors such as limited training data (due to the complexity of curating experimental specimens or high-fidelity simulations) and environmental volatility of many industrial settings (resulting in the corruption of on-the-fly measurements). Thus, an end-to-end machine learning (ML) workflow for acoustic weld inspection in realistic (i.e., industrial) settings has remained an elusive goal. This work addresses the challenges of data curation and signal corruption by proposing workflow consisting of a reduced-order modeling scheme, diffusion based distribution alignment, and U-Net-based segmentation and inversion. A reduced-order Helmholtz model based on Lamb wave theory is used to generate a comprehensive dataset over varying weld heterogeneity and crack defects. The relatively inexpensive low-order solutions provide a robust training dataset for inversion models which are refined through a transfer learning stage using a limited set of full 3D elastodynamic simulations. To handle out-of-distribution (OOD) real-world measurements with varying and unpredictable noise distributions, i.e., Laser Doppler Vibrometry scans, guided diffusion produces in-distribution representations of OOD experimental LDV scans which are subsequently processed by the inversion models. This integrated framework provides an end-to-end solution for automated weld inspection on real data.
@@ 2   Acoustic Model and Forward Problem
Curating a sufficiently large and diverse datasets that faithfully represent the dominant characteristics of the forward problem (i.e., steady-state wave propagation over weldlines) is a necessary precursor for any ML-based inversion task. This is a challenge for weld inspection [38] since high-fidelity simulations capturing all conceivable phenomena (e.g., anisotropy, nonlinearity, etc) are notoriously expensive and challenging to perform [39]. While some researchers have utilized generative techniques to address this [27, 30], such methods still require a large dataset to train augmentative models or risk biasing generations toward a narrow training distribution, limiting dataset diversity. [40]. Hence, we propose a hierarchy of forward model complexities as an alternative. First, a high-fidelity 3D elastodynamic Navier-Lamé (NL) model that captures the dominant scattering physics, but at a high computational expense. Next, a low-order 2D effective medium (EM) model utilizing Lamb wave theory to construct 2D scalar-valued models that capture the dominant wave behavior of the 3D elastodynamic process at a reduced computational cost. The NL model, being far more expensive to query, serves both to calibrate EM solutions and to provide a low-volume dataset (roughly 10% of the entire training set) of full fidelity solutions for ML model tuning.
@@ 2.1   Elastodynamic NL Model
The NL model is constructed as follows. Let $\Omega=[0,L]\times[0,W]\times[0,H]\in\mathbb{R}^{3}$ with external boundary $\partial\Omega$ be the 3D scattering domain with L,W,H>0L,\ W,\ H>0 being the length, width, and height. Let $\Omega_{w}\subset\Omega$ be a welded region along a 2D path $\gamma(s),s=(x(s),y(s))$ with narrow voids representative of cracks in weldlines, $\Gamma_{c}\subset\Omega_{w}$. In some simulations, perfectly matched layers (PMLs) are introduced along $\partial\Omega$, denoted as $\Omega_{\mathsf{PML}}$. Time-harmonic tractions are applied in $\Gamma_{t}\subset\partial\Omega$ denoted as $\bm{f}(\bm{x})=-F(x,y)\exp(i\omega t)\bm{e}_{z}$, with $F(x,y)$ being the amplitude profile on the surface $z=h=H/2$, with $\bm{x}=\begin{bmatrix}x&y&z\end{bmatrix}^{\intercal}$ being the position vector, $\omega$ the frequency, and $i$ the imaginary unit. Defining $\Omega_{b}:=\Omega\setminus\Omega_{w}$ as the bulk material region, the interface $\Gamma_{wb}:=\partial\Omega_{w}\cap\partial\Omega_{b}$ is taken as the boundary between weld and bulk-material such that $\overline{\Omega_{w}}\cup\overline{\Omega_{b}}=\overline{\Omega}$. We consider constant material properties within $\Omega_{b}$ and spatially varying properties in $\Omega_{w}$.
@@
The spatial variations of $E(\bm{x})$ are designed to emulate weldline behavior per the baseline material stiffness $E_{0}$. Namely, we consider a composition of several sub-functions: a nominal weldline function $W(\bm{x})$, a bead function ${B}(\bm{x})$ and a pseudo-randomly structured variation function ${V}(\bm{x})$. $W(\bm{x})$ enforces a baseline stiffness reduction (common in weldlines) about a weld radius $r_{w}$, ${B}(\bm{x})$ applies a pseudo-periodic beadline profile to stiffness reduction line (Fig 1(a)), and ${V}(\bm{x})$ takes the form of a smoothly varying Gaussian random field so as to emulate heterogeneity in weldline strength. Altogether, $E(\bm{x})$ is given as,
@@
For NL simulations, we utilize quadratic Lagrange tetrahedral elements using the gmsh-4.3.1 API for meshing, FEnICsx-0.6.0 for system matrix assembly, and PETSc-3.18.5 as the solver engine. Further details on the FE implementation of Eq (1) are given in B. Even with efficient numerical solvers, there is a high memory requirement associated with the 3D vector-valued NL problem on a high-fidelity mesh (Fig 1(c)). This is particularly the case for (relatively) high frequency simulations, whereby at least 6 elements per wavelength are required for quadratic elements; this is further exacerbated by the complexity of welded domains as sufficient mesh refinement is required near the weld line and crack to account for the rapidly varying material properties. For example, a baseline simulation with W=L=12W=L=12", H=0.25H=0.25", and $\omega=2\pi\times 225000$ rad/s necessitates millions of degrees of freedom. Therefore, a lighter simulation procedure is desired; we describe a proposed alternative the subsequent section.
@@ 2.2   Effective Medium Model
Figure 3: Confirmation of the Lamb-wave assumptions. (a) A full D continuum for computing NL solutions and reduced 2D continuum for generating EM solutions. (b) The resulting out-of-plane displacement fields at $z=h$ with their filtered (c) $(A,0)$ and (d) $(S,0)$ components showing qualitative agreements. (e) The spectra of each wavefield show aggregated energy along pre-computed Lamb wave mode numbers (dashed-lines), confirming the alignment of the high-fidelity problem with Lamb theory.
@@ 2.5   Dataset Design
Simulation datasets were generated to ensure sufficient diversity of key domain parameters in both training and testing. Namely, the weld width $L_{w}$, weld angle $\theta_{w}$, force location $\bm{x}_{f}$, nominal weld stiffness reduction $W_{0}$, and relative weld variation amplitude $V_{0}$, are drawn form normal distributions—this emulates deviations from a nominal welding and inspection processes due to natural process variability from nominal welding and measurement schemes. The crack length $L_{c}$ and crack depth $d_{c}$ are both drawn from uniform distributions—this emulated deviations in crack nature which are not suspected to be biased toward any nominal length or depth. Finally, we consider problem domains both with and without cracks, with the probability of a crack occurring denoted as $\mathcal{P}_{c}$ being sampled from a Bernoulli—this prevents biasing the proposed data-driven inversion framework from always predicting cracks when none are present. These parameters are varied per the distributions defined in Table 3 for each problem domain discussed in section 2.4 with the exception of $\theta_{w}$ which is fixed at 0 for the periodic pipe problem to ensure continuity for the periodic boundary conditions. The distributions of Table 3 ensure that the models are trained on a diversity of meta-parameters emulating real-world variations. For instance, scanning fixtures such as scanning laser doppler vibrometers (LDVs) may not always be perfectly oriented, and hence the weld may appear shifted or tilted relative to the scan. Moreover, the placement of transducers is subject to change, so $\bm{x}_{f}$ in the simulated models should emulate this. Lastly, weldlines may vary in width and relative strength, and these nominal parameters must be diversely sampled as well to mitigate bias in training.
@@ 3   Machine Learning Framework
The machine learning methodology addressed two tasks: wavefield inversion and distribution alignment. The former is a set of inverse models that are responsible for ingesting wavefield data, either directly from simulation test sets, real-world experiments, or from the output of a conditional generation. The latter serves to draw out-of-distribution (OOD) experimental samples into the training distribution, i.e., the EM and NL simulations, via conditional generative modeling.
@@ 3.1   Inversion Models
The baseline architectures of $\mathsf{U}_{{\text{inv}}}$ and $\mathsf{U}_{{\text{seg}}}$ are both comprised of 5 encoding and decoding blocks, with convolution block attention (CBAM) layers being applied after each convolution block. Aside from this, $\mathsf{U}_{{\text{seg}}}$ incorporates batch normalization to stabilize training, as well as a multi-axis feature extraction (MAFE) module following CBAM to improve spatial contextualizing in the segmentation task, whereas $\mathsf{U}_{{\text{inv}}}$ incorporates Fourier Neural Operator (FNO) modules between skipped connections to capture global dependencies efficiently in the frequency domain, facilitating accurate modeling of complex acoustical fields. The workflow draws inspiration from the recent results of Ref [43], which demonstrated the advantages of FNO, CBAM, and MAFE modules for forward modeling of discontinuous parametric domains. The details of each module are detailed below.
@@ 3.1.3 Multi-Axis Feature Extraction
The Multi-Axis Feature Extraction (MAFE) module is designed to capture both global and local dependencies along spatial axes in order to strengthen contextual representation for segmentation; similar methodologies has proven effective in various image segmentation tasks [48, 49]. Herein, we utilize the same MAFE module described Ref [43]. Given an input feature map ${\mathsf{F}}\in\mathbb{R}^{B\times C\times H\times W}$, where $B$, $C$, $H$, and $W$ denote batch, channel, height, and width, respectively, MAFE applies axis-wise gated MLPs along the horizontal ($f_{m_{x}}$) and vertical ($f_{m_{i}}$) directions after normalization, producing a global response, $\mathsf{g}=f_{m_{x}}({\mathsf{F}})+f_{m_{y}}({\mathsf{F}})$. To complement this with local information, the feature map is partitioned into non-overlapping blocks of size $\tfrac{H}{G}\times\tfrac{W}{G}$ ($G=2$ herein), unfolded into patches, and processed by local axis MLPs $f^{l}_{x},f^{l}y$. This yields the local response
@@ 3.1.4 Inverse Model Training Objectives
Three competing objectives guide the weld parameter regression model. The first is an MSE loss between the true and predicted parameter map on $\bar{\bm{x}}$, denoted $\phi(\bar{\bm{x}})$ and scaled by $\mathsf{w}_{1}$—its purpose is self-evident. Second, a gradient loss encourages smoothness in non-welded (flat) regions, which is scaled by $\mathsf{w}_{2}$. Finally, a focal loss scaled by $\mathsf{w}_{3}$ exploits local and spatially compact variations. The composite loss is thus,
@@ 3.2   Super-resolution Conditional Diffusion Model
The inversion framework outlines in section 3.1 is exclusively trained on simulated datasets. Ergo, generalizing its performance in a zero-shot fashion to experimental measurements—for which data is too scarce to train with—must be achieved. Experimental LDV meta-parameters alter the underlying distribution of noise and artifacts in unpredictable ways, making it difficult to conventionally denoise experimental measurements or to synthetically noise training data to represent the diversity of noise distributions that may be present in live deployment. Moreover, even the high-fidelity NL solutions are simplified in the sense that they neglect nonlinearity and anisotrpy, and thus miss the full complexity of real-world measurements, meaning that even clean signals require distribution re-alignment when processing experimental data. Therefore, we propose an alternative to conventional data pre-processing—a generative process that is trained to produce wavefield representations from the training (simulation) distribution based on sparse or corrupted wavefield representations from physical measurements. Specially, we consider a conditioned denoising diffusion probabilistic model (DDPM), detailed herein.
@@ 4   Results
We utilize numerical and physical experiments to evaluate the enrichment training scheme, the model performance on hold-out NL simulations, and the adaptability to real-world data. We first demonstrate the EM solutions (2.2) provide meaningful data enrichment for learning the inverse of high-fidelity NL models (2.1).This is followed by qualitative and quantitative performance metrics for each of the problem classes described in section 2.4. Lastly, the performance of the DDPM-driven distribution alignment described by section 3.2 and downstream inverse task is shown for the experimental data described in section 2.6.
@@ 4.2   Experimental Results
The inference results on direct measurements are given by Fig 13(a), showing the real-component of $\bm{\varphi}^{\text{exp}}$ for each scan speed as well as the model’s crack and weld stiffness predictions on $\mathsf{X}^{\text{exp}}$. The models struggle to achieve either weld characterization or crack identification on the direct measurements. Although the large crack of the experimental sample is identified for all scan speeds, it is accompanied by many false positives. As the scan rate increases, the smaller crack is lost, and the predicted crack maps seem to lose all meaning with respect to the ground truth depicted in Fig 5. Moreover, the parameter maps produced by $\mathsf{U}_{\rm inv}$ are difficult to interpret, with discontinuities and a non-physically noisy boundaries shown. Hence, as expected, the model fails to generalize to OOD experimental measurements directly.
Fig 13(b) depicts the inverse same results applied to conditionally generated fields, $\bm{\varphi}^{\text{exp}}_{0}$. Now, in contrast, exceptional performance is recovered for the crack detection problem, with the predicted mask $\hat{\mathcal{C}}\left(\mathsf{X}^{\text{exp}}_{0}\right)$ indicating cracks in the same two regions flagged by the calibrated ground-truth mask of Fig 5. Moreover, the inference for the weld properties are now consistent with the expected result: nominal bulk properties in the unwelded domain, and a noticeable (15-20 percent) stiffness reduction with in the weldline itself. Hence, the proposed framework distribution alignment of OOD experimental measurements to the known training distribution (based on 3D elastodynamic modeling) is shown effective for applying the inverse model to real-world measurements.